Chloe invested​ $500 by purchasing 50 call options for XYZ stock at ​$85 per share. At its highest​ point, the stock reached $ 97 per share. If Chloe had exercised her option at that​ time, what would her return have​ been? 

A. 500%
B. 970%
C. 120%
D. 20%
E. 83.3%
Answer #1

Invested Amount = $500

Profit on Call Option = (97 - 85)50 = $600

Rate of Return = 600/500

Rate of Return = 120%